OpenStudy (anonymous):

-16t^2 + 18t + 7 = 0 **Solve by using the quadratic equation. After the equation is solved, to the nearest tenth of a second, when will the ball have a height of 10ft?

OpenStudy (anonymous):

t = [ -18 +/- sqrt(18^2 - 4*-16*7) ] /2*-16 t = [ -18 +/- sqrt(324 + 448) ]/-32 t = [ -18 +/- 27.8 ] /-32 t = 1.43 seconds is positive time answer -------------------------------------------- h = Hi + Vi t - 16 t^2 if h, the height, and Hi, the initial height, are feet and g is 32 ft/s^2 (obsolete units but that is what you are using) in this case Hi seems to be 7 and Vi = 18 10 = 7 + 18 t - 16 t^2 16 t^2 - 18 t + 3 = 0 t = [ 18 +/- sqrt (324 - 192) ]/32 t = [ 18 +/- 11.5 ] /32 t = .20 seconds on the way up or t = .92 seconds on the way down